OpenStudy (anonymous):

Simplify. x^2 ÷ x^-3

OpenStudy (neer2890):

\[\frac{ x ^{2} }{ x ^{-3} }=x ^{2}*x ^{3}\]

OpenStudy (neer2890):

when base is same ,powers are added and base remains same.

OpenStudy (anonymous):

x^-6?

OpenStudy (neer2890):

x^2 multiplied by x^3 base is same(x) add powers of x(2+3) and put it on same base =x^5